OBSERVATION OF PURE ROTATIONAL SPECTRA OF trans-, and cis-HOCO/DOCO

\underline{TAKAHIRO OYAMA}, WATARU FUNATO, YOSHIHIRO SUMIYOSHI and YASUKI ENDO, Department of Basic Science, Graduate School of Arts and Sciences, The University of Tokyo, Komaba, Meguro-ku, Tokyo, 153-8902, Japan.

Rotational spectra of trans-, cis-HOCO and DOCO have been observed using an Fourier transform microwave(FTMW) spectrometer and the FTMW-mmw-DR technique. The cis-conformer was observed for the first time in the gas phase. The HOCO radical was produced in a supersonic jet by discharging a gas mixture of CO and H_{2}O diluted in Ar. The molecular constants including the hyperfine coupling constants have been precisely determined. The determined r_{0} structure of the two conformers are consistent with those of ab initio calculations. The fermi constants of the two conformers indicate that the spin densities on the hydrogen nuclei are different between the two conformers.